# Grey-body factors of higher dimensional regular black holes in quasi-topological theories

### v1

@@ III WKB approach for the calculation of grey-body factors
$\Gamma_{\ell}(\omega)\equiv|T|^{2}=1-|R|^{2}.$ (10)
Applying the WKB approximation to Eq. (7) allows one to compute the scattering coefficients in terms of the WKB phase through the relation
$\Gamma_{\ell}(\omega)=\frac{1}{1+e^{2i\pi\mathcal{K}}},$ (11)
@@
Following the original formulation of Schutz and Will Schutz and Will (1985), we employ the WKB expansion corrected to higher orders derived in Iyer and Will (1987); Konoplya (2003); Matyjasek and Opala (2017); Matyjasek and Telecka (2019). The resulting expression reads
$\mathcal{K}\;=\;i\frac{\omega^{2}-V_{0}}{\sqrt{-2V_{0}^{\prime\prime}}}-\sum_{k=2}^{N}\Lambda_{k}.$ (12)
where $V_{0}$ denotes the value of the effective potential at its maximum $r=r_{max}$, and $V_{0}^{\prime\prime}$ is the second derivative of the potential with respect to the tortoise coordinate evaluated at the same point. The correction terms $\Lambda_{k}$ depend on higher-order derivatives of the potential at $r_{max}$. The above WKB approach was broadly used for finding quasinormal modes and grey-body factors of black holes Konoplya and Abdalla (2005); Konoplya (2020); Mahdavian Yekta et al. (2021); Konoplya and Zinhailo (2020b); Matyjasek (2021); Konoplya (2021); Konoplya and Zhidenko (2023); Bolokhov (2024b); Dubinsky and Zinhailo (2025); Bolokhov (2024c); Skvortsova (2024); Malik (2025d); Miyachi et al. (2025); Hamil and Lütfüoğlu (2025); Pedrotti and Calzà (2025); Lütfüoğlu (2025c); Antonelli et al. (2025); Zhang et al. (2025).
@@
Figure 1: Difference between WKB approximations of various orders and the accurate grey-body data for electromagnetic perturbations of the $D=5$ black hole $(f)$ for $\ell=1$ and $\alpha=0.3r_{0}^{2}$: $V_{S}$ (8a) (top panel) and $V_{V}$ (8b) (bottom panel). Accurate data are taken from Ref. Konoplya and Zhidenko (2025b).
@@
An alternative method for estimating the grey-body factors was proposed in Ref. Konoplya and Zhidenko (2024c, 2025c). This approach expresses the transmission coefficient in terms of the two dominant quasinormal modes, $\omega_{0}$ and $\omega_{1}$. The quasinormal mode spectrum for the present class of spacetimes was studied using the WKB method in Ref. Arbelaez (2025). The resulting approximate expression for the grey-body factor reads
$\Gamma_{\ell}(\omega)\approx\left[1+\exp\left(\frac{2\pi\bigl(\omega^{2}-\mathrm{Re}(\omega_{0})^{2}\bigr)}{4\,\mathrm{Re}(\omega_{0})\,\mathrm{Im}(\omega_{0})}\right)\right]^{-1}+\mathcal{O}(\ell^{-1}),$ (13)
which provides accurate results for low multipole numbers and becomes exact in the eikonal limit $\ell\to\infty$. The above correspondence has been applied and tested in a number of recent publications Skvortsova (2025c); Lütfüoğlu (2025d, 2026); Lütfüoğlu et al. (2025c); Huang and Li (2025); Liang et al. (2026), showing as a rule a good concordance for higher multipole numbers, provided the effective potential has a single peak. For double well potentials the correspondence is usually broken Konoplya and Stashko (2025). Notice that the grey-body factors are usually more stable to small deformations of the background than quasinormal modes Oshita et al. (2024); Konoplya and Pappas (2025).
IV Energy emission rate
Figure 2: Transmission coefficients for electromagnetic perturbations of the $D=5$ black hole $(f)$: $V_{S}$ (8a) (top panel) and $V_{V}$ (8b) (bottom panel). The curves show the dependence on the parameter $\alpha=0,0.1,\ldots,0.6,0.66$, increasing from darker to lighter tones.
@@ IV Energy emission rate
$\frac{\partial^{2}E}{\partial\omega\partial t}=\frac{1}{2\pi}\sum_{\ell}{N_{\ell}\Gamma_{\ell}(\omega)\frac{\omega}{\exp{(\omega/T_{H})-1}}},$ (14)
@@
$T_{H}=\frac{f^{\prime}(r_{0})}{4\pi},$ (15)
and $N_{\ell}$ are the multiplicity factors accounting for the degeneracy of electromagnetic modes in $D$-dimensional spacetime. For scalar-type ($V_{S}$) and vector-type ($V_{V}$) perturbations, these factors are given by Kanti (2004)
$N_{\ell}^{(S)}$ $=$ $\frac{(2\ell+D-3)(\ell+D-4)!}{\ell!(D-3)!},$ (16)
@@
Figure 3: Hawking temperature with $D=5$ and $r_{0}=1$ for various regular black holes: black line $(a)$, grey line $(b)$, blue line $(c)$, red line $(d)$, pink line $(e)$, and orange line $(f)$.
@@
Figure 4: Energy emission rates for the electromagnetic perturbations of the $D=5$ black hole $(f)$: $V_{S}$ (8a) (top panel) and $V_{V}$ (8b) (bottom panel), for different values of the parameter $\alpha$.
@@
0 $1.13194\times 10^{-3}$ $7.14534\times 10^{-4}$
0.1 $6.13596\times 10^{-4}$ $3.51646\times 10^{-4}$
0.2 $2.72738\times 10^{-4}$ $1.37896\times 10^{-4}$
0.3 $8.75558\times 10^{-5}$ $3.74379\times 10^{-5}$
0.4 $1.59197\times 10^{-5}$ $5.51799\times 10^{-6}$
0.5 $1.39108\times 10^{-6}$ $4.59524\times 10^{-7}$
0.6 $1.26795\times 10^{-7}$ $4.34701\times 10^{-8}$
Table 2: Energy emitted per unit time for the electromagnetic perturbations ($D=5$, $r_{0}=1$): $V_{S}$ (8a) and $V_{V}$ (8b) for different values of $\alpha$ for black hole $(f)$.
@@
$\frac{dM}{dt}=-\frac{dE}{dt}\;\simeq\;-\,\frac{1}{r_{0}^{2}}\sum_{i}\Xi_{i}(r_{0}),$ (17)
@@
The strong suppression of the emission rate as $\alpha$ approaches its upper bound suggests the exponential dependence governed primarily by the Hawking temperature given by Eq. (15). We fit the numerical data using the following function:
$\Xi_{i}(r_{0})\;=\;A_{i}\,\exp\!\left(-\,\frac{1+a_{i}\dfrac{\alpha}{r_{0}^{2}}+b_{i}\dfrac{\alpha^{2}}{r_{0}^{4}}+c_{i}\dfrac{\alpha^{3}}{r_{0}^{6}}}{r_{0}\,T_{H}(r_{0})}\right),$ (18)
which provides an accurate description of both scalar-type and vector-type electromagnetic emission for all black-hole configurations considered. The numerical values of the coefficients $a_{i}$, $b_{i}$, $c_{i}$, and $A_{i}$ are reported in Tables (3-4). As $\alpha$ increases, the dominant contribution to the exponent arises from the inverse Hawking temperature $1/T_{H}$, since $T_{H}$ decreases more rapidly than the geometric corrections encoded in the higher-order terms.
To express the evaporation dynamics in terms of the geometric size of the black hole, we use Eq. (5) to write the mass as a function of the horizon radius,
$M(r_{0})=\frac{(D-2)\Omega_{D-2}}{16\pi G_{D}}\mu(r_{0}),$ (19)
where, from Eq. (4),
@@
Differentiating Eq. (19) and using Eqs. (3), (4), and (15), after some algebra we find
@@
Combining this expression with the energy-emission law (17), we obtain the differential equation for the horizon radius,
$\frac{dr_{0}}{dt}=\frac{8G_{D}(1+2\pi r_{0}T_{H}(r_{0}))\sum_{i}\Xi_{i}(r_{0})}{(D^{2}-3D+2)\Omega_{D-2}T_{H}(r_{0})r_{0}^{D+1}h\left[\frac{1}{r_{0}^{2}}\right]}.$ (20)
@@
Assuming that the energy-emission rates for all perturbation types and channels exhibit the same qualitative dependence on the coupling parameter $\alpha$ as found for the electromagnetic field, one can infer the general features of the evaporation process. For $\alpha=0$, the Hawking temperature is $T_{H}(r_{0})\propto r_{0}^{-1}$, the quantities $\Xi_{i}(r_{0})$ remain constant, and the black hole evaporates completely in a finite time. In contrast, for $\alpha>0$, Eq. (20) admits a stable equilibrium point characterized by $T_{H}(r_{0})=0$, corresponding to the maximal allowed value of $\alpha/r_{0}^{2}$. The horizon radius then approaches this remnant configuration only asymptotically, as the emission rate becomes exponentially suppressed.
V Conclusions
@@
In this work, we investigated the propagation of electromagnetic fields in a class of higher–curvature regular black hole spacetimes arising from quasi-topological gravity. These geometries, generated by infinite–curvature corrections, provide singularity-free black hole solutions while preserving a relatively simple analytic structure for static, spherically symmetric backgrounds. Our analysis focused on the impact of such corrections on grey-body factors.
@@
Our results show that, for all the regular black holes under consideration, the potential barrier becomes higher and broader, leading to a suppression of low-frequency transmission. Using the computed grey-body factors, we evaluated the energy-emission rates for electromagnetic radiation. We found that increasing of the coupling parameter $\alpha$ to a substantial suppression of the total radiated power. The emission spectra shift toward lower frequencies and diminish by several orders of magnitude as the coupling approaches its maximal allowed value. This behavior is well captured by an exponential fit dominated by the inverse Hawking temperature.
@@
We also considered an effective evolution equation for the horizon radius during evaporation. Under the assumption that other perturbative channels exhibit a similar dependence on $\alpha$, our analysis suggests a qualitative change in the evaporation process. While the Schwarzschild-like case $\alpha=0$ leads to complete evaporation in a finite time, the presence of higher–curvature corrections results in an asymptotic approach to a remnant configuration characterized by vanishing Hawking temperature and exponentially suppressed radiation.
@@
These results indicate that regular black holes may exhibit distinctive observational signatures, potentially allowing them to be distinguished from the singular black hole solutions of General Relativity. The suppression of grey-body factors and the associated modifications of the Hawking emission spectrum provide concrete imprints of the underlying higher–curvature structure of the spacetime. While the present analysis focused on electromagnetic fields as probes of these effects, graviton emission is expected to provide the dominant contribution in higher-dimensional spacetimes. A detailed study of gravitational perturbations and their corresponding emission channels therefore constitutes a natural extension of this work.

### v2

@@ 3 WKB approach for the calculation of grey-body factors
$\Gamma_{\ell}(\omega)\equiv|T|^{2}=1-|R|^{2}.$ (3.2)
Applying the WKB approximation to Eq. (2.7) allows one to compute the scattering coefficients in terms of the WKB phase through the relation
$\Gamma_{\ell}(\omega)=\frac{1}{1+e^{2i\pi\mathcal{K}}},$ (3.3)
@@
Following the original formulation of Schutz and Will [107], we employ the WKB expansion corrected to higher orders derived in [48, 67, 93, 94]. The resulting expression reads
$\mathcal{K}\;=\;i\frac{\omega^{2}-V_{0}}{\sqrt{-2V_{0}^{\prime\prime}}}-\sum_{k=2}^{N}\Lambda_{k}.$ (3.4)
where $V_{0}$ denotes the value of the effective potential at its maximum $r=r_{max}$, and $V_{0}^{\prime\prime}$ is the second derivative of the potential with respect to the tortoise coordinate evaluated at the same point. The correction terms $\Lambda_{k}$ depend on higher-order derivatives of the potential at $r_{max}$. The above WKB approach was broadly used for finding quasinormal modes and grey-body factors of black holes [52, 68, 88, 73, 95, 69, 59, 15, 33, 13, 109, 92, 97, 43, 104, 81, 4, 116].
@@
Figure 1: Difference between WKB approximations of various orders and the accurate grey-body data for electromagnetic perturbations of the $D=5$ black hole $(f)$ for $\ell=1$ and $\alpha=0.3r_{0}^{2}$: $V_{S}$ (2.8a) (left) and $V_{V}$ (2.8b) (right). Accurate data are taken from ref. [75].
@@
An alternative method for estimating the grey-body factors was proposed in Ref. [60, 63]. This approach expresses the transmission coefficient in terms of the two dominant quasinormal modes, $\omega_{0}$ and $\omega_{1}$. The quasinormal mode spectrum for the present class of spacetimes was studied using the WKB method in Ref. [5]. The resulting approximate expression for the grey-body factor reads
$\Gamma_{\ell}(\omega)\approx\left[1+\exp\left(\frac{2\pi\bigl(\omega^{2}-\mathrm{Re}(\omega_{0})^{2}\bigr)}{4\,\mathrm{Re}(\omega_{0})\,\mathrm{Im}(\omega_{0})}\right)\right]^{-1}+\mathcal{O}(\ell^{-1}),$ (3.5)
which provides accurate results for low multipole numbers and becomes exact in the eikonal limit $\ell\to\infty$. The above correspondence has been applied and tested in a number of recent publications [112, 82, 83, 86, 47, 77], showing as a rule a good concordance for higher multipole numbers, provided the effective potential has a single peak. For double well potentials the correspondence is usually broken [54]. Notice that the grey-body factors are usually more stable to small deformations of the background than quasinormal modes [101, 72].
4 Energy emission rate
Figure 2: Transmission coefficients for electromagnetic perturbations of the $D=5$ black hole $(f)$: $V_{S}$ (2.8a) (top panel) and $V_{V}$ (2.8b) (bottom panel). The curves show the dependence on the parameter $\alpha=0,0.1,\ldots,0.6,0.66$, increasing from darker to lighter tones.
@@ 4 Energy emission rate
$\frac{\partial^{2}E}{\partial\omega\partial t}=\frac{1}{2\pi}\sum_{\ell}{N_{\ell}\Gamma_{\ell}(\omega)\frac{\omega}{\exp{(\omega/T_{H})-1}}},$ (4.1)
@@
$T_{H}=\frac{f^{\prime}(r_{0})}{4\pi},$ (4.2)
and $N_{\ell}$ are the multiplicity factors accounting for the degeneracy of electromagnetic modes in $D$-dimensional spacetime. For scalar-type ($V_{S}$) and vector-type ($V_{V}$) perturbations, these factors are given by [50]
$N_{\ell}^{(S)}$ $=$ $\frac{(2\ell+D-3)(\ell+D-4)!}{\ell!(D-3)!},$ (4.3)
@@
Figure 3: Hawking temperature for $D=5$ and $D=8$ with $r_{0}=1$ for various regular black holes: black line $(a)$, grey line $(b)$, blue line $(c)$, red line $(d)$, pink line $(e)$, and orange line $(f)$.
@@
Figure 4: Energy emission rates for the electromagnetic perturbations of the $D=5$ black hole $(f)$: $V_{S}$ (2.8a) (left) and $V_{V}$ (2.8b) (right), for different values of the parameter $\alpha$.
@@
0 $1.131\,94\text{\times}{10}^{-3}$ $7.145\,34\text{\times}{10}^{-4}$
[…]
0.2 2.727 $27\,38\text{\times}{10}^{-4}$ $1.378\,96\text{\times}{10}^{-4}$
0.3 $8.755\,58\text{\times}{10}^{-5}$ $3.743\,79\text{\times}{10}^{-5}$
0.4 $1.591\,97\text{\times}{10}^{-5}$ $5.517\,99\text{\times}{10}^{-6}$
0.5 $1.391\,08\text{\times}{10}^{-6}$ $4.595\,24\text{\times}{10}^{-7}$
0.6 $1.267\,95\text{\times}{10}^{-7}$ $4.347\,01\text{\times}{10}^{-8}$
Table 3: Energy emitted per unit time for the electromagnetic perturbations ($D=5$, $r_{0}=1$): $V_{S}$ (2.8a) and $V_{V}$ (2.8b) for different values of $\alpha$ for black hole $(f)$.
@@
$\frac{dM}{dt}=-\frac{dE}{dt}\;\simeq\;-\,\frac{1}{r_{0}^{2}}\sum_{i}\Xi_{i}(r_{0}),$ (4.4)
@@
The strong suppression of the emission rate as $\alpha$ approaches its upper bound suggests the exponential dependence governed primarily by the Hawking temperature given by Eq. (4.2). We fit the numerical data using the following function:
$\Xi_{i}(r_{0})\;=\;A_{i}\,\exp\!\left(-\,\frac{1+a_{i}\dfrac{\alpha}{r_{0}^{2}}+b_{i}\dfrac{\alpha^{2}}{r_{0}^{4}}+c_{i}\dfrac{\alpha^{3}}{r_{0}^{6}}}{r_{0}\,T_{H}(r_{0})}\right),$ (4.5)
which provides an accurate description of both scalar-type and vector-type electromagnetic emission for all black-hole configurations considered. The numerical values of the coefficients $a_{i}$, $b_{i}$, $c_{i}$, and $A_{i}$ are reported in Tables (4-5). As $\alpha$ increases, the dominant contribution to the exponent arises from the inverse Hawking temperature $1/T_{H}$, since $T_{H}$ decreases more rapidly than the geometric corrections encoded in higher-order terms.
To express the evaporation dynamics in terms of the geometric size of the black hole, we use Eq. (2.5) to write the mass as a function of the horizon radius,
$M(r_{0})=\frac{(D-2)\Omega_{D-2}}{16\pi G_{D}}\mu(r_{0}),$ (4.6)
where, from Eq. (2.4),
@@
Differentiating Eq. (4.6) and using Eqs. (2.3), (2.4), and (4.2), after some algebra we find
@@
Combining this expression with the law on energy-emissions (4.4), we obtain the differential equation for the horizon radius,
$\frac{dr_{0}}{dt}=\frac{8G_{D}(1+2\pi r_{0}T_{H}(r_{0}))\sum_{i}\Xi_{i}(r_{0})}{(D^{2}-3D+2)\Omega_{D-2}T_{H}(r_{0})r_{0}^{D+1}h\left[\frac{1}{r_{0}^{2}}\right]}.$ (4.7)
@@
Assuming that the energy-emission rates for all types of perturbation and channels exhibit the same qualitative dependence on the coupling parameter $\alpha$ as found for the electromagnetic field, one can infer the general features of the evaporation process. For $\alpha=0$, the Hawking temperature is $T_{H}(r_{0})\propto r_{0}^{-1}$, the quantities $\Xi_{i}(r_{0})$ remain constant, and the black hole evaporates completely in a finite time. In contrast, for $\alpha>0$, Eq. (4.7) admits a stable equilibrium point characterized by $T_{H}(r_{0})=0$, corresponding to the maximal allowed value of $\alpha/r_{0}^{2}$. The horizon radius then approaches this remnant configuration only asymptotically, as the emission rate becomes exponentially suppressed [37].
5 Conclusions
@@
In this work, we investigated the propagation of electromagnetic fields in a class of higher–curvature regular black hole spacetimes arising from quasi-topological gravity. These geometries, generated by infinite–curvature corrections, provide singularity-free black hole solutions while preserving a relatively simple analytic structure for static, spherically symmetric backgrounds. Our analysis focused on the impact of such corrections on the grey-body factors.
@@
Our results show that, for all the regular black holes under consideration, the potential barrier becomes higher and broader, leading to a suppression of low-frequency transmission. Using the computed grey-body factors, we evaluated the energy emission rates for electromagnetic radiation. We found that increasing the coupling parameter $\alpha$ leads to a substantial suppression of the total radiated power. The emission spectra shift toward lower frequencies and decrease by several orders of magnitude as the coupling approaches its maximal allowed value. This behavior is well captured by an exponential fit dominated by the inverse Hawking temperature.
@@
We also considered an effective evolution equation for the horizon radius during evaporation. Under the assumption that other perturbative channels exhibit a similar dependence on $\alpha$, our analysis suggests a qualitative modification of the evaporation process. While the Schwarzschild-like case $\alpha=0$ leads to complete evaporation in finite time, the presence of higher–curvature corrections results in an asymptotic approach to a remnant configuration characterized by a vanishing Hawking temperature and exponentially suppressed radiation.
@@
These results indicate that regular black holes may exhibit distinctive observational signatures, potentially allowing them to be distinguished from the singular black hole solutions of General Relativity. The suppression of grey-body factors and the associated modifications of the Hawking emission spectrum provide concrete imprints of the underlying higher–curvature structure of spacetime. Although the present analysis focused on electromagnetic fields as probes of these effects, graviton emission is expected to provide the dominant contribution in higher-dimensional spacetimes. A detailed study of gravitational perturbations and their corresponding emission channels therefore constitutes a natural extension of this work.